OpenStudy (anonymous):

Find the slope of the line that contains each pair of points (5,3) (-2, 0.5) Find the slope of the line that contains each pair of points (5,3) (-2, 0.5) @Mathematics

OpenStudy (mimi_x3):

Use the gradient formula y2-y1 / x2-x1 x1 = 5 , y1=3 ; x2= -2 , y2=0.5

OpenStudy (anonymous):

Slope = Rise / Run Slope = Y1 - Y2 / X1 - X2 (5,3) y1 = 3, x1 =5 (-2, 0.5) y2 = 0.5, x2 = -2 Slope = (3 - 0.5 )/(5 - (-2)) Slope = 2.5/7 or 0.357142857